The Briggs and Stratton Commercial Division designs and manufacturers small engines for golf turf maintenance equipment. A robotics-based testing system with support equipment will ensure that their new signature guarantee program entitled "Always Insta-Start" does indeed work for every engine produced.

Pull System Push System
First cost of equipment $-2,000,000 $-2,700,000
AOC per Year $-700,000 $-440,000
Salvage Value $110,000 $70,000
Estimated Life 8 years 8 years

Compare the annual worth of the two systems at MARR = 9% per year. Select the better system.

The   (Click to select)  push system  pull system  is determined to be the better system.
Answer #1

Given,

MARR = 9% per year

Pull System:

First cost of equipment = -$2,000,000

From the compound interest table, we obtain (A/P, 9%, 8) = 0.1807

The annual worth of the first cost = -$2,000,000 x (A/P, 9%, 8) = -$2,000,000 x 0.1807 = -$361,400

Salvage value at the end of 8 years = $110,000

From the compound interest table, we btain (A/F, 9%, 8) = 0.0907

The annual worth of the salvage value = $110,000 x (A/F, 9%, 8) = $110,000 x 0.0907 = $9,977

Annual Operating Costs = -$700,000

Total annual worth of the pull system = AW of the first cost + AOC + AW of the Salvage Value = -$361,400 - $700,000 + $9,977 = -$1,051,423

Push System:

First cost of equipment = -$2,700,000

From the compound interest table, we obtain (A/P, 9%, 8) = 0.1807

The annual worth of the first cost = -$2,700,000 x (A/P, 9%, 8) = -$2,700,000 x 0.1807 = -$487,890

Salvage value at the end of 8 years = $70,000

From the compound interest table, we btain (A/F, 9%, 8) = 0.0907

The annual worth of the salvage value = $70,000 x (A/F, 9%, 8) = $70,000 x 0.0907 = $6,349

Annual Operating Costs = -$440,000

Total annual worth of the pull system = AW of the first cost + AOC + AW of the Salvage Value = -487,890 - $440,000 + $6,349 = -$921,541

The annual worth of the pull system is less than that of push system (Or, alternately, it can be said that the cost incurred is lower for push system) . Therefore, push system is the better one.